Troy Engines, Ltd., manufactures a variety of engines for use in heavy equipment. The company has always produced all of theComplete this question by entering your answers in the tabs below. Required 1 Required 2 | Required 3 Required 4 Assuming theRequired 1 Required 2 Required 3 Required 4 Should the outside suppliers offer be accepted? Yes ONO Required 1 Required 3 >Required 1 Required 2 Required 3 Required 4 Suppose that if the carburetors were purchased, Troy Engines, Ltd., could use theRequired 1 Required 2 Required 3 Required 4 Given the new assumption in requirement 3, should the outside suppliers offer be
Answer #1

1) Differential analysis

Make Buy
Direct material 330000
Direct labor 176000
Variable overhead 66000
Fixed overhead 22000
Purchase cost (22000*35) 770000
Total relevant cost 594000 770000

Financial (disadvantage) = 594000-770000 = -176000

2) No

3) Differential analysis

Make Buy
Direct material 330000
Direct labor 176000
Variable overhead 66000
Fixed overhead 22000
Opportunity Cost 220000
Purchase cost (22000*35) 770000
Total relevant cost 814000 770000

Financial advantage = 814000-770000 = 44000

4) Yes